Session: Nov 2001 (session 3)
Wavelength: 6.0 cm
To give feedback on this experiment, click
Shanghai (Huang Xinyong at Thu Nov 15 04:02:41 MET 2001) Success
The scan No0042 was lost because of operating with trouble.
Medicina (Tonino Pisanu at Thu Nov 15 08:55:16 MET 2001) Success
Many sources where too low in elevation.
Effelsberg (A. Kraus at Thu Nov 15 09:02:28 MET 2001) Minor failures
Several sources were below our elevation limit. Hardware and software
problems lead to the loss of data between UT0917-0922, UT1607-1619,
Urumqi (Liu Xiang at Thu Nov 15 09:45:21 MET 2001) Minor failures
the power was down from UT04:38:52 to 05:29:26, see logfile,
and the first scan lost.
Onsala (John Conway at Thu Nov 15 10:49:22 MET 2001) Minor failures
The last couple of hours were missed due to high winds.
Westerbork (Tony Foley at Thu Nov 15 13:36:46 MET 2001) Minor failures
13 of the 14 telescopes in array (RTB had pointing control software
problems).Many scans at start and all scans after 18:06UT out
of hour angle range. Stopped schedule at 18:07.
Westerbork (Tony Foley at Thu Nov 15 13:37:17 MET 2001) Minor failures
Missed scan of 4C39.25 at 03:15UT -bug in scheduling!
13:04 - 13:10 OCS software
13:30 - 13:50 DCB delay control stopped
16:40 - 17:15 OCSmanager
17:30 - 17:50 HA stuck at max
Westerbork (Tony Foley at Thu Nov 15 13:38:04 MET 2001) Minor failures
first tape sent to JIVE for fringe test with Eff. Maybe a problem
with the lost scan of 4C39.25
Jodrell_Bank (Alastair Gunn at Thu Nov 15 14:44:00 MET 2001) Minor failures
Telescope taken offline at 17:07 to drive out of azimuth
limit. Back on source at 17:21. Tsys values wrong for this
experiment because diode was incorrectly switched.
Torun (K.Borkowski at Fri Nov 16 17:20:39 MET 2001) Failure
Did not participate because of continued terminal problems.
Urumqi (Liu Xiang at Tue Nov 20 03:37:05 MET 2001) Minor failures
observer told me that the date might be wrong,it was found after
observation of ed019a.